OpenStudy (anonymous):

the sum of two consecutive integers is at least 14. find the least possible pair of integers

OpenStudy (anonymous):

pls show work !!! thanks

OpenStudy (anonymous):

Let the two consequtive integers be p and (p+1) then, p+(P+1)=14 2p=13 p=13/2 Hence the first no is 13/2 and second no is (13/2)+1=15/2 Therefore the two consequtive integers are 13/2 and 15/2